OpenStudy (anonymous):

A student is opens a math book to two facing pages. The product of the page numbers is 156. Find the page numbers?

OpenStudy (saifoo.khan):

156/2

OpenStudy (anonymous):

I divided 156 in half and got 78 but now I am not sure if the pages are 77 and 78 or 78 and 79 I am confused a little.

OpenStudy (anonymous):

let one page no. =x the other will be x+1 now x(x+1)=156 solve!!

OpenStudy (anonymous):

you want to solve x(x+1) = 156

OpenStudy (dumbcow):

no you are looking for 2 consecutive numbers that multiply to 156 n*(n+1) = 156 n^2 + n - 156 = 0 (n-12)(n+13) = 0 n must be positive n= 12

OpenStudy (anonymous):

The product of the page numbers is 156, not the sum. Lets call the page numbers n and n+1, so\[n(n+1) = 156 \Rightarrow n^2 + n - 156 = 0 \Rightarrow n = 12, -13\]Since the page number cannot be negative, \[n = 12\]and the page numbers are n and n+1, which are 12 and 13.

OpenStudy (anonymous):

So 12 and 13 are the page numbers

OpenStudy (anonymous):

If so then how did I come up with the answer that I go?

OpenStudy (anonymous):

you got 78 and 79, but their product isnt 156, so those cant be the answer. the product is 6162